OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find [f(2)-g(1)]². A. 1 B. 2 C. 9

OpenStudy (anonymous):

5-4=1 a

OpenStudy (anonymous):

the way you'd write this out is: \[f(2)= (2)^{2} +1\] and then \[g(1)=3(1) + 1\] when you subtract g(1) from f(2), you would get \[((2)^{2}+1)-(3*(1)+1)= 5-4 = 1\] And then, since you want to SQUARE this answer, you would get \[(1)^{2}= 1\] Therefore, your answer is A! Good luck!